A manager is trying to decide whether to purchase a certain part or to have it produced internally. Internal production could use either of two processes. One would entail a variable cost of $17 per unit and an annual fixed cost of $200,000; the other would entail a variable cost of $14 per unit and an annual fixed cost of $240,000. Three vendors are willing to provide the part. Vendor A has a price of $20 per unit for any volume up to its maximum capacity of 30,000 units. Vendor B has a price of $22 per unit for demand less than 1,000 units, and $18 per unit for larger quantities. Vendor C offers a price of $21 per unit for the first 1,000 units, and $19 per unit for additional units.

  

a.

If the manager anticipates an annual volume of 10,000 units, which alternative would be best from a cost standpoint? For 20,000 units, which alternative would be best? (Omit the "$" sign in your response.)

   

  TC for 10,000 units TC for 20,000 units
  Int. 1: $ Int. 1: $
  Int. 2: $ Int. 2: $
  Vend A $ Vend A $
  Vend B $ Vend B $
  Vend C $ Vend C $
(Click to select)Vendor AInt.1Int.2Vendor BVendor C is the best from a cost standpoint. (Click to select)Vendor BVendor AVendor CInt.1Int.2 is the best from a cost standpoint.
Answer #1

For 10,000 Units

Total cost = Fixed cost + Variable cost

TC for Int1: 200,000 + 17*10000 = 200,000 + 170,000 = 370000

TC for Int 1: 240,000 +14*10000 = 240000+140000 = 380000

TC for VendA : 20*10000 = 200000

TC for VendB: 18*10000 = 180,000

TC for VendC : 21*1000 + 19*9000 = 192000

Clearly, Vend B is the best for overall cost stand point

For 20,000 Units

Total cost = Fixed cost + Variable cost

TC for Int1: 200,000 + 17*20000 = 200,000 + 340,000 = 540000

TC for Int 1: 240,000 +14*20000 = 240000+280000 = 520000

TC for VendA : 20*20000 = 400000

TC for VendB: 18*20000 = 360,000

TC for VendC : 21*1000 + 19*19000 = 382000

Clearly, Vend B is the best for overall cost stand point